OpenStudy (anonymous):

2(x-5)-x(x+1)=x(2-x) help.

OpenStudy (anonymous):

Simplifying 2(x + -5) + -1x(x + 1) = x(2 + -1x) Reorder the terms: 2(-5 + x) + -1x(x + 1) = x(2 + -1x) (-5 * 2 + x * 2) + -1x(x + 1) = x(2 + -1x) (-10 + 2x) + -1x(x + 1) = x(2 + -1x) Reorder the terms: -10 + 2x + -1x(1 + x) = x(2 + -1x) -10 + 2x + (1 * -1x + x * -1x) = x(2 + -1x) -10 + 2x + (-1x + -1x2) = x(2 + -1x) Combine like terms: 2x + -1x = 1x -10 + 1x + -1x2 = x(2 + -1x) -10 + 1x + -1x2 = (2 * x + -1x * x) -10 + 1x + -1x2 = (2x + -1x2) Add 'x2' to each side of the equation. -10 + 1x + -1x2 + x2 = 2x + -1x2 + x2 Combine like terms: -1x2 + x2 = 0 -10 + 1x + 0 = 2x + -1x2 + x2 -10 + 1x = 2x + -1x2 + x2 Combine like terms: -1x2 + x2 = 0 -10 + 1x = 2x + 0 -10 + 1x = 2x Solving -10 + 1x = 2x Move all terms containing x to the left, all other terms to the right. Add '-2x' to each side of the equation. -10 + 1x + -2x = 2x + -2x Combine like terms: 1x + -2x = -1x -10 + -1x = 2x + -2x Combine like terms: 2x + -2x = 0 -10 + -1x = 0 Add '10' to each side of the equation. -10 + 10 + -1x = 0 + 10 Combine like terms: -10 + 10 = 0 0 + -1x = 0 + 10 -1x = 0 + 10 Combine like terms: 0 + 10 = 10 -1x = 10 Divide each side by '-1'. x = -10 Simplifying x = -10

OpenStudy (anonymous):

Did this help ?

OpenStudy (anonymous):

yes thank you :)

OpenStudy (anonymous):

You're welcome :)